OpenStudy (anonymous):

Calculus Question: A rancher plans to make four identical and adjacent rectangular pens, each with an area of 16m^2. What are the dimensions of each pen that minimize the amount of fence that must be used?

OpenStudy (campbell_st):

ok... will start with area of 1 pen A = length x width or A = lw so 16 = lw so let w = 16/l Perimeter P = 2l + 2w or substituting P = 2l + 16/l you now need to find the 1st derivative dP/dl and then let it equal zero and solve

OpenStudy (anonymous):

Thanks!

OpenStudy (campbell_st):

I thin you'll find each pen needs to be a 4x4 square... so the total fencing is 48m